OpenStudy (anonymous):

a. Find the surface area of the cube. b. Find the volume of the cube. 4x-5 is one side length

OpenStudy (anonymous):

surface area of a cube with side a=6a^2 and volume=a^3

OpenStudy (anonymous):

now put in the a as 4x-5

OpenStudy (anonymous):

and see if u get d answer?

OpenStudy (anonymous):

ans.a) 6a^2=6(4x-5)=24x-30 b)volume=(4x-5)^3

OpenStudy (anonymous):

jeyaram its 6a squared, u used sumthing else;)